CourtneyWatson:

I need help with my math homework plz it is due in the morning it is about Special Case Solutions I will post the questions down below:

CourtneyWatson:

How many solutions does the equation: 3x+15=2x+10+x+5 have? The equation has_______________ solutions.

CourtneyWatson:

@Hoodmemes can you or someone else please help?

CourtneyWatson:

3x+15=2x+10+x+5 3x+15=_____x +______ 3x=______+15=3x-________+15 ______=______

CourtneyWatson:

Please someone help

CourtneyWatson:

If the value of the x is negative would the equation still be true? Explain

CourtneyWatson:

@mxddi3

SelfishQueen:

did u get help?

Hobii:

@IamAku

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) CourtneyWatson 3x+15=2x+10+x+5 3x+15=_____x +______ 3x=______+15=3x-________+15 ______=______ \(\color{#0cbb34}{\text{End of Quote}}\) Ok so you would solve \(3x+15=2x+10+x+5\) by first adding parenthesis to combine the like terms Once you add the parenthesis it would look like this \(3x+15=(2x+x)+(10+5)\) So combining the like terms, \((2x+x=3x)\) because if there is \(x\) with no coefficient then it is equal to \(1x\), and 2+1=3, so 3x then 10+5 which is 15 So \(3x+15=(2x+x)+(10+5)=\) \(3x+15=3x+15\) So then the second thing we do is get the \(3x+\) Change the plus `(+)` to a minus `(–)` So then we \(–3x\) (because we changed to to a –) to both sides of the equation So \(3x+15−3x=3x+15−3x\) That solved would equal \(15=15\) So the last step, \(–15\) to both sides of the equation, so \(-15\) to both sides of \(\bf 15=15\) \(15−15=15−15\) so \(\sf ?=?\) Using this you should be able to fill in the blansk

supie:

Blanks*

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @CourtneyWatson How many solutions does the equation: 3x+15=2x+10+x+5 have? The equation has_______________ solutions. \(\color{#0cbb34}{\text{End of Quote}}\) @supie here the question not was answered pls. what is the right answer ? what mean this case when you get simplified every terms or how is in this case when the left side is equal the right side ?

jhonyy9:

how many solutions has this equation ?

supie:

It would be \(0=0\) so that means there will only be one solution right?

supie:

brb afk

jhonyy9:

no you get that for indifferent value of x the left side will be equal with the right side so what result from this ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok pls remember in this case there are infinity solutions bc. you can give indifferent values to x ALWAYS the left side will be equal right side so infinity solutions \(\color{#0cbb34}{\text{End of Quote}}\)

jhonyy9:

@Laylalyssa

supie:

Ohh ok Thanks @jhonyy9

jhonyy9:

np yw